### v1

# Spectroscopy of 4He at 0.25 ppt Uncertainty and Improved Alpha–Helion Charge-Radius Difference Determination

Abstract
High-precision spectroscopy of simple atomic systems can be used to advance the theory of atomic energy levels but can also serve as a sensitive probe of nuclear charge radii. For this last purpose, we report an improved measurement of the transition frequency in 4He with 48 Hz uncertainty (0.25 ppt), using a Bose–Einstein condensed sample confined in a magic-wavelength optical dipole trap. A systematic Doppler shift from condensate motion is suppressed by time-resolved ion detection, and the transition frequency is calibrated via a White Rabbit link to a remote active hydrogen maser clock. Combined with previous 3He measurements and improved theory, we obtain the most precise determination to date of the charge-radius difference between the alpha and helion particles of of . This is consistent with other recent determinations and confirms that the current discrepancy between QED theory and experimentally observed ionization energies of excited states in helium is not apparent in the isotope shift.
Precision measurements of the energy structure of simple atomic and molecular systems have provided some of the most stringent tests of quantum electrodynamics (QED) and, in general, of the standard model; see e.g. [1, 2, 3, 4]. The fundamental constants required for theory calculations can be obtained by combining theory with high-precision experimental data [5, 4]. By comparing the results based on different systems, a valuable consistency check can be done in the comparison between experiment and theory. In this way, spectroscopy of electronic [6, 7, 8] and muonic hydrogen [9] has led to adjustments of the Rydberg constant and the proton radius [5].
Helium, as the simplest multi-electron system after hydrogen, serves as an important platform for testing QED and two-electron effects. Combined with the experimental benefits of its laser-coolable and efficiently detectable state, helium provides a unique platform for testing fundamental physics [10, 11].
QED calculations of triplet helium (both electron spins parallel) have advanced to a level that includes corrections up to [12]. This means that finite-size effects from the nucleus start to play an important role. However, significant discrepancies exist between these theoretical calculations and the respective measurements of the transition [13], transition [14], and ionization energy [15, 16]. So far, it is unknown where these discrepancies stem from. This makes it impossible to determine the absolute size of the helium nucleus directly from spectroscopic measurements.
Instead, the isotope shift on a transition in and can be measured. The cancellation in the isotope shift of mass-independent terms allows subkilohertz accuracy of the theoretical calculations, so that the differential nuclear charge radius squared between the alpha and helion particles can be determined from it [17]. Comparing measurements of from different systems provides a broad consistency check, testing not only the various experiments but also the different calculations required to extract .
This is illustrated by the latest developments: recently, in [18] a large discrepancy [19] between various measurements [20, 21, 22] of from the was resolved by identifying a systematic postselection shift. Also, the measurement from the was improved [10], and the previous difference of [23] with the in helium was explained by a systematic shift due to the interplay between the Fermi gas and the optical trapping potential. A comparison of obtained with electronic (regular) helium with a measurement from muonic helium ions revealed a surprising 3.6 difference [24]. This inspired more advanced calculations [25, 26] of the hyperfine mixing in revealing a shift of kHz in the isotope shift calculation for the for a point-like nucleus [26], reducing the difference of compared to muonic helium ions to 1.2[10]. Interestingly, has also been determined from the isotope shift of the ionization energies of the state [27]. Although there is a discrepancy between theory and the calculated ionization energy, the determined agrees with measurements from both muonic and electronic helium, suggesting that the discrepancy is not caused by finite size effects.
In this work, we present the most precise frequency measurement in helium on the transition in , leading to an improved determination of the charge radius difference between the helion and the alpha particle. The transition is observed in a Bose-Einstein Condensate (BEC), confined in a magic-wavelength optical dipole trap. Calibration of the transition frequency is performed via a White Rabbit fiber link referenced to a remote active hydrogen maser. We developed a method to observe the motion of the BEC in the optical trap and methods to reduce this motion and its effects on the calibration. In addition, we determined the singlet-triplet scattering length and the magic wavelength more precisely.
The starting point for our spectroscopy measurements is a BEC of in the metastable state trapped in an optical dipole trap near the magic wavelength of 319.816 nm. A schematic overview of our experimental setup is shown in Figure 1. Details on how we produce quantum degenerate gasses of He* in our experiment can be found in [28, 10, 29].
The dipole trap light is generated by frequency-doubling 640 nm light, produced via sum-frequency generation of amplified 1557 nm spectroscopy light and a separate amplified 1085 nm laser, whose wavelength is stabilized to a precision wavemeter (High Finesse WSU-30) to set the trap wavelength. The power of the dipole trap laser is controlled using an acousto-optic modulator (AOM). This enables control over the number of atoms and the chemical potential of the BEC by ramping down the trap power to expel atoms, followed by ramping it back up to obtain the desired trap depth.
The spectroscopy laser (bandwidth ) is transfer-locked to an ultrastable laser via an optical frequency comb [30]. The frequency comb and counters used for metrology are referenced via a 110 km long White Rabbit fiber link to a hydrogen maser located at VSL, the national metrology institute of the Netherlands. This link reaches a stability in the low after of averaging, with a noise floor of .
The atoms in the BEC are all prepared from the start in one of the two spin-stretched states to suppress collision-induced Penning ionization. We then expose the BEC to spectroscopy light for 100 ms. When a triplet atom is excited to a singlet state, it collides with one of the many surrounding triplet atoms and Penning ionizes, producing a He+ or He ion. These ions are detected using a negatively biased microchannel plate (MCP), and the detection times of these ions are recorded with a time-to-digital converter (TDC). The number of counts as a function of laser frequency is our spectroscopy signal. In the 100 ms preceding the excitation, background counts (from ionization of background gas [28] and the scattered light from the dipole trap beams) are recorded. After exposure to the spectroscopy light, the atoms are released and fall under gravity onto a second MCP positioned below the optical dipole trap. By fitting the time-of-flight traces recorded with this MCP, we determine the chemical potential and the atom number of the BEC.
To measure a spectrum, we repeat the above procedure for each spectroscopy laser frequency. We apply a magnetic field of 0.75 G, which is actively stabilized to cancel out changes in the ambient magnetic field between measurements. We alternate between exciting from the and the states to the 2 state and take the average to find the magnetic field-free transition frequency. Figure 2 shows an example of a measured spectrum. The observed linewidth of kHz is attributed to a combination of mean-field broadening, magnetic noise at 50 Hz from the AC-line, the AC magnetic field generated by the turbomolecular pump used to evacuate the science chamber, and residual Doppler broadening from axial trap oscillations of the BEC.
Figure 2 shows an example of a histogram of the ion (signal) arrival times over a 100 ms spectroscopy exposure time. Peaks are observed when the velocity of the oscillating BEC induces a Doppler shift such that the atoms become resonant with the frequency of the spectroscopy laser. In [28], axial trap oscillations of the BEC, induced by beam pointing noise of the dipole trap beams, were identified as the dominant broadening mechanism. By increasing the mechanical stability of the setup, we have suppressed these oscillations due to pointing noise. However, we still observe oscillations induced by the switching of the magnetic trap and bias coils. Since the oscillations have a well-defined phase relationship with the experimental sequence, they can lead to a systematic frequency shift. We implemented two methods to cancel shifts. (i) In post-processing, we set the spectroscopy exposure time equal to the axial trapping period, which is measured daily. This ensures that each measurement spans a complete oscillation cycle and that the net frequency shift averages to zero. (ii) We record two spectra, one for each of the two counter-propagating spectroscopy beam directions. This results in two data sets measured from opposite directions. By taking the average of both sides, we can suppress any residual Doppler shift.
The three systematic shifts that need to be extrapolated to zero are the ac Stark shift from the spectroscopy laser, the ac Stark shift from the dipole trap, and the mean-field shift arising from interactions between the atoms in the BEC. These frequency shifts scale linearly with the spectroscopy laser power, dipole trap power, and chemical potential, respectively. We measure these effects by varying the three parameters independently and fit a three-dimensional weighted linear regression model to the data to find the transition frequency.
There is a 50 ms delay between the end of the spectroscopy interaction and the release of the BEC from the trap. During this time, we turn off the bias coils, and if the atoms are in the state, we transfer them to the state for the measurement of the chemical potential. During this time, the chemical potential decays exponentially due to off-resonant scattering by the dipole trap light. We corrected the measured chemical potential for this decay. When all is taken into account, the extrapolation for a single measurement set looks like the typical example shown in Figure 3a-d.
The results of all nine measurement sets (based on 698 spectra) are shown in Figure 3e-f and were recorded between 28 November 2024 and 26 January 2025. The weighted averages of the two sides differ by . To further suppress Doppler shifts, we combine both sides by taking their arithmetic mean. We estimate that the residual Doppler shift is less than (see Supplementary Material).
The measured transition frequency (see Table 1) is affected by several systematic shifts and uncertainties that need to be taken into account. The most dominant, but easily calculable, correction arises from the recoil shift due to the absorption of a 1557 nm photon. The next leading effect originates from magnetic field gradients caused by magnetic materials near the setup. If the states are spatially displaced within such a gradient, they experience different magnetic fields, and averaging the transition frequencies obtained from them does not completely cancel the Zeeman shift. To quantify this effect in our experiment, we compared the Zeeman shift experienced by the states by driving RF spin flips between both states, comparing the resonance conditions for atoms first prepared in to those that started in . We found a difference of Hz, which is taken into account in Table 1. The uncertainty in the one-body lifetime, which we use to correct the measured chemical potentials, leads to an uncertainty of in the extrapolation to the unperturbed transition frequency; this includes the decay of the chemical potential during excitation. To check for any nonlinearities in the extrapolation of the chemical potential, we repeated the fit using only data points with chemical potentials greater than and less than . No systematic deviation from the results presented here was found. Detailed derivations of all systematic shifts given in Table 1 can be found in the supplementary material [31].
| Value/Correction | Value (Hz) | Uncertainty (Hz) |
| Measured frequency | 27 | |
| Recoil shift | 0.01 | |
| Magnetic field gradient | -6 | 22 |
| Decay | 0 | 33 |
| Residual Doppler shift | 0 | 3 |
| Blackbody radiation | 0.3 | |
| DC-Stark shift ion MCP | 0 | |
| Clock | 0 | |
| Gravitational redshift | 0.03 | |
| Quantum interference | 0.1 | |
| Second-order Zeeman | - | |
| Second-order Doppler | - | |
| Total | 48 | |
| Rengelink et al.[28] | 200 | |
| van Rooij et al.[23] |
After applying all corrections, we find the transition frequency for the to be . Our measurement is in excellent agreement with the previous best result by [28], but it is four times more precise.
For each measurement set, we changed the dipole trap wavelength and determined the corresponding differential ac Stark shift slope. To extract the nearest magic wavelength (where the differential ac Stark shift is zero), we use the (differential) polarizability curve from [32] to fit the measured slopes as a function of the optical dipole trap wavelength. This is shown in Figure 4. The magic wavelength we determined in this manner is equal to 319.81602(4) nm. Our measurement agrees with the previous result reported by [28], but it is four times more precise. It also agrees within 2 with the most accurate theoretical prediction to date of from [33], based on a relativistic full-configuration interaction calculation.
Due to the difference in scattering lengths between the triplet and singlet states, atoms in the singlet state experience a different mean-field potential than those in the triplet state. The resulting frequency shift depends on the density. The inhomogeneous density distribution of the BEC, therefore, leads to broadening of the spectral line. However, due to other broadening mechanisms, this meanfield lineshape is not directly observable. We observe only the average shift as we vary the chemical potential. In terms of the and scattering lengths, this shift is given by [28]:
is the slope from the chemical potential regression. From the average of all measurement sets, including uncertainties for the decay of the chemical potential during spectroscopy, we find . Using the measured value from [34] for the triplet-triplet scattering length we find . This result differs by 3.3 from the result of [28] . We attribute this difference to an underestimation of the uncertainty in that work. Our measurement still differs significantly from the best calculation of [35], based on ab initio calculations of the involved molecular potentials [36]. Details regarding the systematic effects on the magic wavelength and the triplet–singlet scattering length are discussed in [31].
Combining our result for the transition frequency with the measurement in [10], we can determine the isotope shift on this transition. To extract the charge radius difference (squared) from the isotope shift we use theory from [26, 37], which has been improved to include the second order QED nuclear recoil correction, thereby reducing the theoretical uncertainty in the isotope shift calculation for a point nucleus to 107 Hz. Using that theory result and the procedure outlined in [37], we determine an improved value for the differential nuclear charge radius squared between the helion and the alpha particle of . Our value is plotted in Figure 5, along with determinations from other experiments. Our value confirms our previous result [10, 26], but the uncertainty is 30% smaller.
In conclusion, we have measured the transition in with 48 Hz accuracy, corresponding to a relative accuracy of . This constitutes the most precise optical frequency measurement in helium to date. It was achieved with a novel method to suppress systematic Doppler shifts arising from BEC oscillations and by calibrating our spectroscopy laser to a remote active hydrogen maser via a White Rabbit fiber link. Our improved determinations of the magic wavelength and the triplet–singlet scattering length provide stringent tests of the current theory of the helium atom. The reported charge radius difference is the most accurate to date and confirms our previous value. It is in agreement within 1.2 with the value based on muonic helium ion spectroscopy [24] and also within 0.8 with the one based on the ionization energy of the 2 state. This last aspect confirms that the isotope shift is not affected by the 9 discrepancy between the theoretical ionization energies of both states and the experimental values [27].
The accuracy of our determination is now dominated by the uncertainty in the theory of the isotope shift (107 Hz [37]) and the accuracy of 170 Hz from the previous spectroscopy measurement [10]. The aim is to improve both to a similar accuracy as we have now demonstrated for 4He, which (providing sufficiently accurate theory will be available) would enable a determination of with an accuracy of .
Acknowledgements.
We thank Joanne Slijper and Rob Kortekaas for technical support, Lex van der Gracht for the construction of the TDC, and Yuri van der Werf, Raphael Jannin, Frank Cozijn, Haoyuan Li, and Maarten Hoogerland for useful discussions. We acknowledge the support of SURF and VSL for providing access to UTC(VSL) via the SURF Time&Frequency Network. K.E. acknowledges support from the Dutch Research Council (NWO) via grant 16MYSTP and 680-91-108.References
- Tiesinga and Mohr [2023] E. Tiesinga and P. J. Mohr, Tests of fundamental physics, in Springer Handbook of Atomic, Molecular, and Optical Physics, edited by G. W. F. Drake (Springer International Publishing, Cham, 2023) pp. 433–447.
- Morgner et al. [2023] J. Morgner, B. Tu, C. M. König, T. Sailer, F. Heisse, H. Bekker, B. Sikora, C. Lyu, V. A. Yerokhin, Z. Harman, J. R. Crespo López-Urrutia, C. H. Keitel, S. Sturm, and K. Blaum, Stringent test of QED with hydrogen-like tin, Nature 622, 53 (2023).
- Adkins et al. [2022] G. Adkins, D. Cassidy, and J. Pérez-Ríos, Precision spectroscopy of positronium: Testing bound-state qed theory and the search for physics beyond the standard model, Physics Reports 975, 1 (2022), precision spectroscopy of positronium: testing bound-state QED theory and the search for physics beyond the Standard Model.
- Delaunay et al. [2023] C. Delaunay, J.-P. Karr, T. Kitahara, J. C. J. Koelemeij, Y. Soreq, and J. Zupan, Self-consistent extraction of spectroscopic bounds on light new physics, Phys. Rev. Lett. 130, 121801 (2023).
- Tiesinga et al. [2021] E. Tiesinga, P. J. Mohr, D. B. Newell, and B. N. Taylor, Codata recommended values of the fundamental physical constants: 2018, Rev. Mod. Phys. 93, 025010 (2021).
- Beyer et al. [2017] A. Beyer, L. Maisenbacher, A. Matveev, R. Pohl, K. Khabarova, A. Grinin, T. Lamour, D. C. Yost, T. W. Hänsch, N. Kolachevsky, and T. Udem, The Rydberg constant and proton size from atomic hydrogen, Science 358, 79 (2017).
- Fleurbaey et al. [2018] H. Fleurbaey, S. Galtier, S. Thomas, M. Bonnaud, L. Julien, F. Biraben, F. Nez, M. Abgrall, and J. Guéna, New measurement of the transition frequency of hydrogen: Contribution to the proton charge radius puzzle, Phys. Rev. Lett. 120, 183001 (2018).
- Bezginov et al. [2019] N. Bezginov, T. Valdez, M. Horbatsch, A. Marsman, A. C. Vutha, and E. A. Hessels, A measurement of the atomic hydrogen lamb shift and the proton charge radius, Science 365, 1007 (2019).
- Pohl et al. [2010] R. Pohl, A. Antognini, F. Nez, F. D. Amaro, F. Biraben, J. M. R. Cardoso, D. S. Covita, A. Dax, S. Dhawan, L. M. P. Fernandes, A. Giesen, T. Graf, T. W. Hänsch, P. Indelicato, L. Julien, C.-Y. Kao, P. Knowles, E.-O. L. Bigot, Y.-W. Liu, J. A. M. Lopes, L. Ludhova, C. M. B. Monteiro, F. Mulhauser, T. Nebel, P. Rabinowitz, J. M. F. dos Santos, L. A. Schaller, K. Schuhmann, C. Schwob, D. Taqqu, J. F. C. A. Veloso, and F. Kottmann, The size of the proton, Nature 466, 213 (2010).
- van der Werf et al. [2025] Y. van der Werf, K. Steinebach, R. Jannin, H. L. Bethlem, and K. S. E. Eikema, Alpha and helion particle charge radius difference determined from quantum-degenerate helium, Science 388, 850 (2025).
- Henson et al. [2022] B. M. Henson, J. A. Ross, K. F. Thomas, C. N. Kuhn, D. K. Shin, S. S. Hodgman, Y.-H. Zhang, L.-Y. Tang, G. W. F. Drake, A. T. Bondy, A. G. Truscott, and K. G. H. Baldwin, Measurement of a helium tune-out frequency: an independent test of quantum electrodynamics, Science 376, 199 (2022).
- Patkóš et al. [2021] V. Patkóš, V. A. Yerokhin, and K. Pachucki, Complete lamb shift of helium triplet states, Phys. Rev. A 103, 042809 (2021).
- Dorrer et al. [1997] C. Dorrer, F. Nez, B. de Beauvoir, L. Julien, and F. Biraben, Accurate measurement of the two-photon transition frequency in helium: New determination of the lamb shift, Phys. Rev. Lett. 78, 3658 (1997).
- Luo et al. [2016] P.-L. Luo, J.-L. Peng, J. Hu, Y. Feng, L.-B. Wang, and J.-T. Shy, Precision frequency measurements of transitions at 588 nm, Phys. Rev. A 94, 062507 (2016).
- Clausen et al. [2021] G. Clausen, P. Jansen, S. Scheidegger, J. A. Agner, H. Schmutz, and F. Merkt, Ionization energy of the metastable state of from rydberg-series extrapolation, Phys. Rev. Lett. 127, 093001 (2021).
- Clausen et al. [2025] G. Clausen, K. Gamlin, J. A. Agner, H. Schmutz, and F. Merkt, Metrology in a two-electron atom: The ionization energy of metastable triplet helium , Phys. Rev. A 111, 012817 (2025).
- Pachucki et al. [2017] K. Pachucki, V. Patkóš, and V. A. Yerokhin, Testing fundamental interactions on the helium atom, Phys. Rev. A 95, 062510 (2017).
- Wen et al. [2025] J.-L. Wen, J.-D. Tang, Y.-N. Lv, Y. R. Sun, C.-L. Zou, J.-F. Dong, and S.-M. Hu, Postselection shifts the transition frequency of helium in an atomic beam, Science Advances 11, eadu9796 (2025).
- Zheng et al. [2017] X. Zheng, Y. R. Sun, J.-J. Chen, W. Jiang, K. Pachucki, and S.-M. Hu, Measurement of the frequency of the transition of , Phys. Rev. Lett. 119, 263002 (2017).
- Pastor et al. [2004] P. C. Pastor, G. Giusfredi, P. D. Natale, G. Hagel, C. de Mauro, and M. Inguscio, Absolute frequency measurements of the atomic helium transitions around 1083 nm, Phys. Rev. Lett. 92, 023001 (2004).
- Cancio Pastor et al. [2012] P. Cancio Pastor, L. Consolino, G. Giusfredi, P. De Natale, M. Inguscio, V. A. Yerokhin, and K. Pachucki, Frequency metrology of helium around 1083 nm and determination of the nuclear charge radius, Phys. Rev. Lett. 108, 143001 (2012).
- Shiner et al. [1995] D. Shiner, R. Dixson, and V. Vedantham, Three-nucleon charge radius: A precise laser determination using 3He, Phys. Rev. Lett. 74, 3553 (1995).
- van Rooij et al. [2011] R. van Rooij, J. S. Borbely, J. Simonet, M. D. Hoogerland, K. S. E. Eikema, R. A. Rozendaal, and W. Vassen, Frequency metrology in quantum degenerate helium: Direct measurement of the transition, Science 333, 196 (2011).
- Schuhmann et al. [2025] K. Schuhmann, L. M. P. Fernandes, F. Nez, M. A. Ahmed, F. D. Amaro, P. Amaro, F. Biraben, T.-L. Chen, D. S. Covita, A. J. Dax, M. Diepold, B. Franke, S. Galtier, A. L. Gouvea, J. Götzfried, T. Graf, T. W. Hänsch, M. Hildebrandt, P. Indelicato, L. Julien, K. Kirch, A. Knecht, F. Kottmann, J. J. Krauth, Y.-W. Liu, J. Machado, C. M. B. Monteiro, F. Mulhauser, B. Naar, T. Nebel, J. M. F. dos Santos, J. P. Santos, C. I. Szabo, D. Taqqu, J. F. C. A. Veloso, A. Voss, B. Weichelt, A. Antognini, R. Pohl, and T. C. Collaboration, The helion charge radius from laser spectroscopy of muonic helium-3 ions, Science 388, 854 (2025).
- Qi et al. [2025] X.-Q. Qi, P.-P. Zhang, Z.-C. Yan, L.-Y. Tang, A.-X. Chen, T.-Y. Shi, and Z.-X. Zhong, Toward resolving the discrepancy in helium-3 and helium-4 nuclear charge radii, Phys. Rev. Res. 7, L022020 (2025).
- Pachucki et al. [2024] K. Pachucki, V. Patkóš, and V. A. Yerokhin, Second-order hyperfine correction to H, D, and energy levels, Phys. Rev. A 110, 062806 (2024).
- Clausen and Merkt [2025] G. Clausen and F. Merkt, Ionization energy of metastable (2 ) and the alpha- and helion-particle charge-radius difference from precision spectroscopy of the rydberg series, Phys. Rev. Lett. 134, 223001 (2025).
- Rengelink et al. [2018] R. J. Rengelink, Y. van der Werf, R. P. M. J. W. Notermans, R. Jannin, K. S. E. Eikema, M. D. Hoogerland, and W. Vassen, Precision spectroscopy of helium in a magic wavelength optical dipole trap, Nature Physics 14, 1132 (2018).
- Jannin et al. [2022] R. Jannin, Y. van der Werf, K. Steinebach, H. L. Bethlem, and K. S. E. Eikema, Pauli blocking of stimulated emission in a degenerate fermi gas, Nature Communications 13, 6479 (2022), published: 29 October 2022.
- Telle et al. [2002] H. Telle, B. Lipphardt, and J. Stenger, Kerr-lens, mode-locked lasers as transfer oscillators for optical frequency measurements, Applied Physics B: Lasers and Optics 74, 1–6 (2002).
- [31] See the supplementary materials.
- Notermans et al. [2014] R. P. M. J. W. Notermans, R. J. Rengelink, K. A. H. van Leeuwen, and W. Vassen, Magic wavelengths for the transition in helium, Phys. Rev. A 90, 052508 (2014).
- Wu et al. [2018] F.-F. Wu, S.-J. Yang, Y.-H. Zhang, J.-Y. Zhang, H.-X. Qiao, T.-Y. Shi, and L.-Y. Tang, Relativistic full-configuration-interaction calculations of magic wavelengths for the 2 transition of helium isotopes, Phys. Rev. A 98, 040501 (2018).
- Moal et al. [2006] S. Moal, M. Portier, J. Kim, J. Dugué, U. D. Rapol, M. Leduc, and C. Cohen-Tannoudji, Accurate determination of the scattering length of metastable helium atoms using dark resonances between atoms and exotic molecules, Phys. Rev. Lett. 96, 023203 (2006).
- Notermans et al. [2016] R. Notermans, R. Rengelink, and W. Vassen, Comparison of spectral linewidths for quantum degenerate bosons and fermions, Physical Review Letters 117, 213001 (2016).
- Müller et al. [1991] M. W. Müller, A. Merz, M. W. Ruf, H. Hotop, W. Meyer, and M. Movre, Experimental and theoretical studies of the Bi-excited collision systems He*(2 3S) He*(2) at thermal and subthermal kinetic energies, Zeitschrift fur Physik D Atoms, Molecules and Clusters 21, 89–112 (1991).
- Pachucki et al. [2025] K. Pachucki, V. Patkóš, and V. A. Yerokhin, QED nuclear recoil effect in helium isotope shift (2025), arXiv:2512.04623 [physics.atom-ph] .
- Huang et al. [2020] Y.-J. Huang, Y.-C. Guan, J.-L. Peng, J.-T. Shy, and L.-B. Wang, Precision laser spectroscopy of the two-photon transition in , Phys. Rev. A 101, 062507 (2020).
- Sick [2014] I. Sick, Zemach moments of and , Phys. Rev. C 90, 064002 (2014).